# MAAT: Mamba Adaptive Anomaly Transformer with association discrepancy for time series

### v1

@@ 3 Methodology
Consider a system that records a sequence of $d$-dimensional measurements at uniform time intervals. The time series data is represented by the set $\{x_{1},x_{2},\dots,x_{N}\}$, where each $x_{t}\in\mathbb{R}^{d}$. The goal is to determine whether an observation $x_{t}$ is anomalous in an unsupervised manner. Effective unsupervised time series anomaly detection relies on learning informative representations and establishing a clear discriminative criterion. The original Anomaly Transformer framework achieved this by learning an association discrepancy that distinguishes standard patterns from anomalies through anomaly attention and a minimax optimization strategy. The Mamba Adaptive Anomaly Transformer (MAAT) builds upon the Anomaly Transformer by introducing computationally efficient Anomaly Sparse Attention, Mamba Blocks, and Gated Skip Connections. This hybrid approach enhances short-range dependency modeling and long-range temporal learning while reducing computational costs. As illustrated in Figure 1(A), Anomaly Sparse Attention replaces dense self-attention with a two-branch mechanism. The Prior-Association Branch encodes expected dependencies using a learnable Gaussian kernel, while the Series-Association Branch applies block-wise sparse attention to capture time series patterns adaptively. These associations guide discrepancy learning, enhancing anomaly detection performance. The Reconstruction Block, shown in Figure 1(B), refines extracted features by stacking MAAT Blocks with LayerNorm and Feed-Forward Networks (FFN). This layered design supports hierarchical feature learning, making the model more robust to diverse time series behaviors. Finally, the MAAT Block, depicted in Figure 1(C), integrates Mamba Blocks within a Gated Attention mechanism, selectively amplifying meaningful signals while filtering out noise, and the skip connection preserves critical information while adapting the model focus. MAAT provides a scalable and expressive framework for detecting complex anomalies across various time series datasets by combining sparse attention, state-space modeling, and gated learning mechanisms with the association discrepancy. Algorithm 1 summarizes the overall computation flow of MAAT, detailing how Anomaly Sparse Attention, Mamba Blocks, and Gated Skip Connections interact to refine the features for anomaly detection.
@@
1:Time series $x\in\mathbb{R}^{B\times L\times D}$
3:Reconstruction $x$, Anomaly scores: $series$, $prior$, $\sigma$
5:$x_{orig}\leftarrow x$triangleright Preserve initial input
6:for $i=1$ to $e\_layers$ do
7:     Sparse Attention Processing:
9:     Mamba Skip Path:
12:     $x_{skip}\leftarrow\text{LayerNorm}(x_{skip})$
13:     Adaptive Gating:
16:     State Update:
17:     $x_{orig}\leftarrow x$triangleright Update input for next layer
18:     Append $series$ to $series\_list$
19:     Append $prior$ to $prior\_list$
20:     Append $\sigma$ to $\sigma\_list$
21:end for
[…]
23:     $x\leftarrow\text{LayerNorm}(x)$
24:end if
25:return $x,series\_list,prior\_list,\sigma\_list$
@@ 3.1.1 Association Discrepancy
The association discrepancy (AssDis) metric quantifies the mismatch between the prior-association ($P$) and the series-association ($S$). The prior-association mechanism models temporal dependencies using a learnable Gaussian kernel, where nearby time steps receive higher weights. The association score between steps $i$ and $j$ in layer $l$ is computed and normalized as:
$P^{l}_{i,j}=\frac{1}{\sqrt{2\pi}\sigma_{i}}\exp\left(-\frac{(j-i)^{2}}{2\sigma_{i}^{2}}\right),\quad P^{l}_{i,:}=\text{Softmax}\left(P^{l}_{i,:}\right),$ (1)
where $\sigma_{i}$ is a learnable scale parameter controlling the temporal influence range.
The series association $S_{i,j}$ represents the observed dependencies between time points using standard self-attention, which computes global attention scores for all pairs of time points. It is defined as:
$S_{i,j}=\frac{\exp\left(\frac{Q_{i}K_{j}^{T}}{\sqrt{d_{\text{model}}}}\right)}{\sum_{k=1}^{T}\exp\left(\frac{Q_{i}K_{k}^{T}}{\sqrt{d_{\text{model}}}}\right)},$ (2)
where $Q_{i}$ and $K_{j}$ are the query and key vectors for the $i$-th and $j$-th time points, respectively, $d_{\text{model}}$ is the dimensionality of the model, used to scale the dot product for numerical stability, and $T$ is the total number of time points in the sequence.
  • 1.
    The prior-association ($P$) captures expected temporal patterns via Gaussian kernels
  • 2.
    The series-association ($S$) encodes actual dependencies through localized attention
  • 3.
    Their discrepancy reveals alignment between expected and observed temporal relationships
The association discrepancy (AssDis) is computed as symmetric KL divergence:
$\text{AssDis}(P,S;X)=\frac{1}{L}\sum_{l=1}^{L}\left[\sum_{i=1}^{N}\text{KL}(P^{l}_{i,:}\parallel S^{l}_{i,:})+\text{KL}(S^{l}_{i,:}\parallel P^{l}_{i,:})\right],$ (3)
  • 1.
    $\text{KL}(P\parallel S)$ measures deviation of observed dependencies from expectations
  • 2.
    Reverse $\text{KL}(S\parallel P)$ prevents mode collapse in divergence measurement
  • 3.
    Symmetric formulation ensures balanced comparison of distributions
3.1.2 Minimax Optimization
@@
In this phase, the parameters responsible for generating $P$ are updated while $S$ is kept fixed (detached). The loss function is defined as:
$\mathcal{L}_{\text{Min}}=\|X-\hat{X}\|_{2}^{2}-\lambda\cdot\text{AssDis}(P,S_{\text{detach}};X),$ (4)
Where:
  • 1.
    $\|X-\hat{X}\|_{2}^{2}$ is the reconstruction loss that ensures the input $X$ is accurately reconstructed.
  • 2.
    $\text{AssDis}(P,S_{\text{detach}};X)$ is the association discrepancy term, which quantifies the dissimilarity between the feature set $P$ and a detached version of $S$. Detaching $S$ prevents its gradients from being updated, providing a stable target for $P$.
  • 3.
    $\lambda$ is a hyperparameter that controls the trade-off between the reconstruction loss and the association discrepancy.
This phase drives $P$ to learn diverse, fine-grained local features while discouraging trivial or collapsed representations.
Maximize Phase:
In this phase, the parameters generating $S$ are updated while $P$ is held fixed (detached). The corresponding loss function is given by:
$\mathcal{L}_{\text{Max}}=\|X-\hat{X}\|_{2}^{2}+\lambda\cdot\text{AssDis}(P_{\text{detach}},S;X),$ (5)
Where:
  • 1.
    The reconstruction loss $\|X-\hat{X}\|_{2}^{2}$ remains in place to ensure accurate reconstruction of the input $X$.
  • 2.
    $\text{AssDis}(P_{\text{detach}},S;X)$ measures the association discrepancy between a detached $P$ and $S$, encouraging $S$ to capture non-local, global dependencies that complement the local features encoded in $P$.
  • 3.
    $\lambda$ again regulates the relative importance of the association discrepancy term.
Adversarial Strategy:
The alternating updates between the two phases create an adversarial learning mechanism:
  • 1.
    In the Minimize Phase, $P$ is encouraged to develop informative and diverse local representations.
  • 2.
    In the Maximize Phase, $S$ is optimized to extract non-local, contextual dependencies.
As introduced in the Anomaly Transformer framework, this minimax strategy ensures that both feature sets capture complementary aspects of the input data, enhancing the overall quality of the learned representations. The hyperparameter $\lambda$ is critical in balancing the network ability to reconstruct the input with the need for diverse and informative feature learning.
@@ 3.1.3 Association-based Anomaly Criterion
This formulation leverages both temporal dependency patterns and reconstruction accuracy for anomaly detection. While improved reconstructions generally reduce the association discrepancy, anomalies may still yield high scores due to significant deviations in either reconstruction error or temporal patterns. The probabilistic weighting ensures that time steps exhibiting both poor reconstruction and abnormal dependencies are emphasized, enhancing detection robustness.
3.2 Sparse Attention for Efficient Sequence Processing
Standard self-attention in Transformer architectures computes interactions between every pair of tokens, resulting in a quadratic computational complexity of $O(N^{2})$. To mitigate this, sparse attention restricts token interactions by applying a sparsity pattern $S$ that selects only a subset of pairs for attention computation [29, 30, 31].
In conventional self-attention, the operation is defined as:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V,$ (7)
@@ 3.2 Sparse Attention for Efficient Sequence Processing
Sparse attention introduces a sparsity mask $S$ to limit interactions:
@@
3.3 Gated Attention Mechanism
The gated attention mechanism enhances standard attention by dynamically adjusting the importance of features through learnable gating. Rather than treating all input features equally, it evaluates their relevance to suppress or amplify contributions, optimizing attention weight distribution. Mathematically, this is expressed as:
$\text{GatedAttention}(x)=\sigma\bigl{(}G(x)\bigr{)}\odot A(x),$ (9)
Where:
  • 1.
    $x$ is the input data,
  • 2.
    $G(x)$ is a neural network-generated gating vector,
  • 3.
    $\sigma$ (sigmoid) maps gating values to $[0,1]$,
  • 4.
    $A(x)$ denotes standard attention weights,
  • 5.
    $\odot$ signifies element-wise multiplication.
The model selectively prioritizes task-critical features while suppressing noise by modulating the attention mechanism with context-sensitive gates. This adaptability improves robustness in scenarios such as anomaly detection, where distinguishing relevant signals from irrelevant background patterns is essential.
@@ 3.4 Mamba State Space Model
3.4.1 State Update Equation
$x_{t+1}=Ax_{t}+Bu_{t}$ (10)
Where:
  • 1.
    $x_{t}$ represents the hidden state at time $t$,
  • 2.
    $A$ is the state transition matrix,
  • 3.
    $B$ is the input matrix,
  • 4.
    $u_{t}$ represents the external input.
The observation equation connects the underlying system dynamics to the measurable signal. It ensures that latent states, which capture long-range dependencies, are projected onto the observed space.
@@ 3.5 Mamba Adaptive Anomaly Transformer Architecture
Figure 1 illustrates the overall structure of the Mamba Adaptive Anomaly Transformer (MAAT), an architecture inspired by the Anomaly Transformer. MAAT integrates a sparse attention mechanism, a selective state-space model known as the Mamba block, and an adaptive reconstruction framework. The sparse attention mechanism processes the input time series data, capturing local dependencies efficiently. The Mamba block is designed to capture long-range dependencies, enhancing the model ability to understand complex temporal patterns. An adaptive gating mechanism then fuses the outputs from the sparse attention and Mamba block, enabling the model to reconstruct the input signal effectively. This combination allows MAAT to detect anomalies by comparing the reconstructed signal with the original input, identifying deviations that may indicate anomalous behavior.
@@ 3.5.1 Anomaly sparse attention
Our anomaly sparse attention module, shown in Figure 1(A), refines the conventional anomaly attention used in the Anomaly Transformer, which relies on self‐attention, by incorporating sparse attention. While the prior-association remains unchanged, modeled using a learnable Gaussian kernel Eq. (1), the series-association is now computed through sparse self-attention Eq. (2), capturing observed dependencies more efficiently.
@@ 3.5.2 MAAT Adaptive Block
The adaptive block shown in Figure 1(C) integrates a state-space model (Mamba block) with a skip connection that retains the original input. The Mamba block is specifically designed to capture long-range dependencies, while the skip connection ensures the preservation of fine details. We achieve a robust intermediate representation by combining these elements and applying layer normalization. This Adaptive Block merges long-range dependencies and local features through a state-space (Mamba) skip path and an adaptive gating mechanism. After performing sparse attention processing, we proceed with the following steps: The Mamba block generates a transformed representation $x_{\text{mamba}}$, which is then combined with the reconstructed input from sparse attention $x_{\text{orig}}$ via a residual connection. This output is normalized as follows:
@@ 4.2 Implmentation
Our study uses the same protocol as the Anomaly Transformer model to assess our approach efficacy. We generate sub-series using a non-overlapping sliding window technique, with most datasets using a fixed window size of 100, while the SMAP dataset uses a size of 105. Anomaly detection involves assigning scores to time points and setting a threshold based on the ”Anomaly Ratio” in Table 1. The batch size varies from 32 to 256 for optimization. We maintain a model dimensionality of 512 for balance and use the ADAM optimizer with early stopping during training. This methodology provides a robust framework for evaluating our approach against existing techniques while ensuring reproducibility and highlighting our improvements.
Table 1: Hyperparameters for MAAT model across Datasets
Dataset Window Size Batch Size $d_{\text{model}}$ Anomaly Ratio (%)
SMD 100 128 512 0.5
MSL 100 256 512 0.85
SMAP 105 128 512 0.5
PSM 100 128 512 1
SWaT 100 256 512 0.5
NIPS-TS-GECCO 100 32 512 0.5
NIPS-TS-SWAN 100 128 512 0.9
@@ 4.3 Evaluation Metrics
Evaluating anomaly detection in time series data requires metrics assessing point-wise performance and capturing anomalies’ temporal continuity. This work uses traditional and specialized point-based metrics to evaluate contiguous anomaly segments.
4.3.1 Point-Based Metrics
4.3.2 Affiliation Metrics: Aff-P and Aff-R
Affiliation Precision (Aff-P):
[…]
Affiliation Recall (Aff-R):
Measures the fraction of true anomalies detected:
[…]
4.3.3 Range-Based Metrics
These assess detection over entire anomaly segments.
Range-based Anomaly Recall (R_A_R):
[…]
Range-based Anomaly Precision (R_A_P):
[…]
4.3.4 Volume-Based Metrics: V_ROC and V_PR
These incorporate anomaly duration into evaluation.
Volume-based ROC (V_ROC):
Adjusts TPR and FPR by anomaly volume:
[…]
The V_ROC curve plots $\text{TPR}_{\text{vol}}$ vs. $\text{FPR}_{\text{vol}}$.
Volume-based Precision-Recall (V_PR):
Defines volume-weighted precision and recall:
In summary, while point-based metrics offer a basic assessment, affiliation, range-based, and volume-based metrics (18, 19, 20, 21) provide a comprehensive evaluation by accounting for temporal structure and anomaly duration [42, 43].
@@ 5.1 Baseline Results
On the SMD dataset, which consists of highly non-stationary multivariate time series with intricate temporal dependencies, our proposed model achieves an F1-score of 92.30%. This performance surpasses the Anomaly Transformer by an absolute margin of +2.18% and outperforms DCdetector by +8.64%. Furthermore, our model demonstrates a notable improvement in Recall, achieving +3.84% higher than the Anomaly Transformer and +13.93% higher than DCdetector, while also exhibiting enhancements in Precision by +0.63% and +3.74%, respectively.
On the MSL dataset, our model attains an F1-score of 95.05%, yielding an absolute improvement of +1.19% over the Anomaly Transformer and +0.32% over DCdetector. In addition, our approach achieves a +2.40% increase in Recall compared to the Anomaly Transformer and +0.95% compared to DCdetector, while demonstrating a +1.14% enhancement in Precision relative to the Anomaly Transformer, with a minor decrease of –0.21% when compared to DCdetector. Similarly, on the SMAP dataset, our method achieves an F1-score of 96.99%, demonstrating an absolute gain of +0.60% over the Anomaly Transformer and +0.93% over DCdetector. Our model also attains a +1.24% improvement in Precision over the Anomaly Transformer and +0.49% over DCdetector. Meanwhile, Recall exhibits a marginal decrease of –0.08% compared to the Anomaly Transformer, but an increase of +1.39% relative to DCdetector. On the SWaT dataset, our model achieves an F1-score of 96.50%, outperforming the Anomaly Transformer by +0.09% and DCdetector by +0.08%. While our model exhibits a –0.28% reduction in Precision compared to the Anomaly Transformer, it surpasses DCdetector by +0.23%. Additionally, our method improves Recall by +0.59% relative to the Anomaly Transformer and by +0.04% compared to DCdetector. Finally, on the PSM dataset, our model achieves an F1-score of 98.32%, demonstrating an improvement of +0.87% over the Anomaly Transformer and +0.50% over DCdetector. Precision is enhanced by +0.35% and +0.27% relative to the Anomaly Transformer and DCdetector, respectively, while Recall is increased by +1.39% and +0.73%, respectively.
Despite these successes, there are instances where MAAT exhibits slightly lower performance compared to specific methods. For example, on the MSL dataset, while MAAT outperforms Anomaly Transformer and closely matches DCdetector, its precision (92.06%) is marginally lower than DCdetector (92.25%), representing a decrease of 0.19%. This minor discrepancy may stem from DCdetector dual-attention contrastive learning approach, which appears to excel in precision for datasets with high variability. Additionally, on the SWaT dataset, MAAT achieves a recall of 100.00%, matching DCdetector, but its precision is slightly lower than DCdetector, representing a decrease of 0.26%. These small trade-offs highlight areas where further optimization could enhance MAAT performance, particularly when precision is prioritized over recall.
@@
MMPCACD 71.20 79.28 75.02 81.42 61.31 69.95 83.22 68.23 74.73 82.52 68.23 74.73 76.26 78.35 77.29
CL-MPPCA 82.36 76.07 79.09 73.71 88.54 80.44 63.16 72.88 67.72 76.78 81.50 79.07 56.02 99.93 71.80
TS-CP2 87.42 66.25 75.38 68.45 68.48 68.42 86.75 83.18 84.95 81.23 74.10 77.50 82.67 78.16 80.35
Deep-SVDD 78.54 79.67 79.10 91.92 76.63 83.26 56.02 69.04 62.40 80.42 84.45 82.39 95.41 86.79 90.73
@@
On the NIPS-TS-GECCO dataset, our model achieves an F1-score of 52.8%, significantly outperforming the Anomaly Transformer (27.0% F1) by an absolute margin of +25.8%. When compared to DCdetector, our model attains an improvement of +6.2% in F1, and further surpasses IForest and OCSVM by +13.7% (52.8% vs. 39.1%) and +23.2%, respectively. In terms of precision, our approach records 42.4%, yielding a +16.7% gain over the Anomaly Transformer (25.7%) and a +4.1% improvement over DCdetector (38.3%). Although IForest exhibits a marginally higher precision of 43.9% (a -1.5% difference relative to our model), this slight trade-off is offset by substantial gains in recall. Specifically, our model achieves a recall of 70.0% compared to 28.5% for the Anomaly Transformer (+41.5%), 59.7% for DCdetector (+10.3%), and 35.3% for IForest (+34.7%), while OCSVM attains a recall of 74.3% (a -4.3% difference relative to our model).
On the NIPS-TS-SWAN dataset, our model exhibits superior performance with the highest precision (95.9%), recall (59.9%), and F1-score (73.8%). In particular, our model outperforms the Anomaly Transformer by +5.2% in precision, +12.5% in the recall, and +11.5% in F1-score. Comparisons with DCdetector reveal more modest gains of +0.4% in precision, +0.3% in recall, and +0.4% in F1-score. In contrast, the improvements over IForest are particularly pronounced, with an absolute increase of +39.0% in precision (95.9% vs. 56.9%) and +15.5% in F1-score. Furthermore, in comparison to OCSVM, our model demonstrates an enhancement of +48.5% in precision and +25.3% in F1-score, along with a +10.1% improvement in recall.
These results highlight the architectural strengths of our approach, especially the Sparse Attention mechanism and Mamba-SSM, which effectively identify subtle anomalies while reducing noise interference. Although minor trade-offs are observed—specifically, a slight reduction in precision on the NIPS-TS-GECCO dataset and a marginal decrease in recall on the NIPS-TS-SWAN dataset—our model consistently balances precision and recall to deliver optimal overall performance.

### v2

@@ 3 Methodology
Consider a system that records a sequence of $d$-dimensional measurements at uniform time intervals. The time series data is represented by the set $\{x_{1},x_{2},\dots,x_{N}\}$, where each $x_{t}\in\mathbb{R}^{d}$. The goal is to determine whether an observation $x_{t}$ is anomalous in an unsupervised manner. Effective unsupervised time series anomaly detection relies on learning informative representations and establishing a clear discriminative criterion. The original Anomaly Transformer framework distinguishes standard patterns from anomalies by learning an association discrepancy via anomaly attention and a minimax optimization strategy. Building on this, the Mamba Adaptive Anomaly Transformer (MAAT) introduces Anomaly Sparse Attention, Mamba Blocks, and Gated Skip Connections, enhancing short-range dependency modeling and long-range temporal learning while reducing computational cost. As illustrated in Figure 1(A), Anomaly Sparse Attention replaces dense self-attention with a two-branch mechanism. The Prior-Association Branch encodes expected dependencies using a learnable Gaussian kernel, while the Series-Association Branch applies block-wise sparse Attention to capture time series patterns adaptively. These associations guide discrepancy learning, enhancing anomaly detection performance. The Reconstruction Block, shown in Figure 1(B), refines extracted features by stacking MAAT Blocks with LayerNorm and Feedforward Networks (FFN). This layered design supports hierarchical feature learning, making the model more robust to diverse time series behaviors. Finally, the MAAT Block, depicted in Figure 1(C), integrates Mamba Blocks within a Gated Attention mechanism, selectively amplifying meaningful signals while filtering out noise, and the skip connection preserves critical information while adapting the model focus. MAAT provides a scalable and expressive framework for detecting complex anomalies across various time series datasets by combining sparse Attention, state-space modeling, and gated learning mechanisms with the association discrepancy.
@@ 3.1.1 Association Discrepancy
As introduced in the Anomaly Transformer framework, this minimax strategy ensures that both feature sets capture complementary aspects of the input data, enhancing the overall quality of the learned representations. The hyperparameter $\lambda$ is critical in balancing the network’s ability to reconstruct the input with the need for diverse and informative feature learning.
@@ 3.1.2 Association-based Anomaly Criterion
This formulation combines temporal patterns and reconstruction accuracy for anomaly detection. While improved reconstructions reduce discrepancy, anomalies with significant deviations in reconstruction or temporal patterns still yield high scores. Probabilistic weighting highlights time steps with poor reconstruction and abnormal dependencies, enhancing robustness.
3.2 Sparse Attention for Efficient Sequence Attentiong
Sparse Attention restricts token interactions by applying a sparsity pattern $S$ that selects only a subset of pairs for attention computation [29, 30, 31]. Sparse attention introduces a sparsity mask $S$ to limit interactions:
@@ 3.2 Sparse Attention for Efficient Sequence Attentiong
The Gated Attention Mechanism enhances standard attention by using a learnable gating process to dynamically adjust the importance of input features. It evaluates the relevance of features to optimize attention weight distribution, represented mathematically as:
$\text{GatedAttention}(x)=\sigma\bigl{(}G(x)\bigr{)}\odot A(x)$
[…]
@@ 3.4 Mamba Adaptive Anomaly Transformer Architecture
Figure 1 illustrates the overall structure of the Mamba Adaptive Anomaly Transformer (MAAT), an architecture inspired by the Anomaly Transformer. MAAT integrates a sparse attention mechanism, a selective state-space model known as the Mamba block, and an adaptive reconstruction framework. The sparse attention mechanism processes the input time series data, capturing local dependencies efficiently. The Mamba block is designed to capture long-range dependencies, enhancing the model’s ability to understand complex temporal patterns. An adaptive gating mechanism then fuses the outpuAttentionhe sparse attention and Mamba block, enabling the model to reconstruct the input signal effectively. This combination allows MAAT to detect anomalies by comparing the reconstructed signal with the original input, identifying deviations that may indicate anomalous behavior.
@@ 3.4.1 Anomaly sparse attention
Our anomaly sparse attention module, shown in Figure 1(A), refines the conventional anomaly attention used in the Anomaly Transformer, which relies on self‐attention and sparse Attention. While the prior-association remains unchanged, modeled using a learnable Gaussian kernelthe series-association is now computed through sparse self-attention, capturing observed dependencies more efficiently.
@@ 3.4.2 MAAT Adaptive Block
The adaptive block shown in Figure 1(C) integrates a state-space model (Mamba block) with a skip connection that retains the original input. The Mamba block is specifically designed to capture long-range dependencies, while the skip connection ensures the preservation of fine details. We achieve a robust intermediate representation by combining these elements and applying layer normalization. This Adaptive Block merges long-range dependencies and local features through a state-space (Mamba) skip path and an adaptive gating mechanism. After performing sparse attention processing, we proceed with the following steps: The Mamba block generates a transformed representation $x_{Mamba{mamba}}$, which is then combined with the reconstructed input from sparse attention $x_{\text{orig}}$ via a residual connection. This output is normalized as follows:
@@ 4.2 Implmentation
Our study follows the Anomaly Transformer model’s protocol to evaluate our approach. We generate a sub-series using a non-overlapping sliding window. Anomaly detection involves scoring time points and setting a threshold based on the Anomaly Ratio detailed in the table. Other parameters, including batch size and model dimensionality, are also specified in Table 7, ensuring robust evaluation and reproducibility of our improvements.
@@ 4.3 Evaluation Metrics
The evaluation of anomaly detection in time series data employs various metrics to assess performance and capture the temporal continuity of anomalies. Point-based metrics, including Precision, Recall, and F1 Score, utilize true positives (TP), false positives (FP), and false negatives (FN) for evaluation[41]. Affiliation metrics, such as Affiliation Precision (Aff-P) and Affiliation Recall (Aff-R), measure the fraction of predicted anomalies within true ranges and the fraction of true anomalies detected, respectively. Range-based metrics like Range-based Anomaly Recall (R_A_R) and Range-based Anomaly Precision (R_A_P) evaluate detection based on the overlap between true and predicted anomaly ranges. Additionally, volume-based metrics, including Volume-based ROC (V_ROC) and Volume-based Precision-Recall (V_PR), adjust true positive rate (TPR) and false positive rate (FPR) by incorporating the volume of anomalies and provide volume-weighted precision and recall. In summary, while point-based metrics offer basic performance insights, affiliation, range-based, and volume-based metrics deliver a more comprehensive evaluation by factoring in the temporal structure and duration of anomalies. In summary, while point-based metrics offer a basic assessment, affiliation, range-based, and volume-based metrics (11, 12, 13, 14) provide a comprehensive evaluation by accounting for temporal structure and anomaly duration [42, 43].
@@ 5.1 Baseline Results
On the SMD dataset, our model outperformed the Anomaly Transformer by an impressive absolute margin of +2.18% and outperformed DCdetector by a substantial +8.64%. Additionally, our model shows a significant improvement in Recall, being +3.84% higher than the Anomaly Transformer and +13.93% higher than DCdetector. It also demonstrates enhancements in Precision of +0.63% and +3.74%, respectively.
On the MSL dataset, our model achieves an absolute improvement of +1.19% in F1-Score over the Anomaly Transformer and +0.32% over DCdetector. Additionally, it shows a +2.40% increase in Recall and a +1.14% enhancement in Precision compared to the former, while experiencing a minor decrease of –0.21% in Precision relative to the latter. Recall also improves by +0.95% compared to DCdetector.
For the SMAP dataset, our method outperformed the Anomaly Transformer by +0.60% and DCdetector by +0.93%. Precision increases by +1.24% and +0.49%, respectively, compared to these models. While Recall shows a slight decline of –0.08% relative to the Anomaly Transformer, it improves by +1.39% over DCdetector.
On the SWaT dataset, our model surpassed the Anomaly Transformer by 0.09% and DCdetector by 0.08% on F1-Score. Although our model shows a –0.28% precision decline compared to the Anomaly Transformer, it exceeds DCdetector by 0.23%. Furthermore, our approach enhances Recall by 0.59% compared to the Anomaly Transformer and by 0.04% concerning the DCdetector.
In the case of the PSM dataset, our model reflected an improvement of 0.87% over the Anomaly Transformer and 0.50% over DCdetector. Furthermore, Precision is enhanced by 0.35% and 0.27% compared to the latter models, respectively, while Recall shows an improvement of 1.39% and 0.73%, respectively.
These minor differences may be attributed to DCdetector’s dual-attention contrastive learning approach, which proves advantageous for datasets characterized by high variability. while on other datasets precision is slightly lower. These small trade-offs underscore opportunities for further optimization to enhance MAAT’s performance, particularly in scenarios where precision is prioritized over recall.
@@
MMPCACD[46] 71.20 79.28 75.02 81.42 61.31 69.95 83.22 68.23 74.73 82.52 68.23 74.73 76.26 78.35 77.29
CL-MPPCA[47] 82.36 76.07 79.09 73.71 88.54 80.44 63.16 72.88 67.72 76.78 81.50 79.07 56.02 99.93 71.80
TS-CP2[48] 87.42 66.25 75.38 68.45 68.48 68.42 86.75 83.18 84.95 81.23 74.10 77.50 82.67 78.16 80.35
Deep-SVDD[49] 78.54 79.67 79.10 91.92 76.63 83.26 56.02 69.04 62.40 80.42 84.45 82.39 95.41 86.79 90.73
@@
On the NIPS-TS-GECCO dataset, Compared to DCdetector, our model shows a notable improvement of 6.2% in F1-score, and it also outperforms IForest and OCSVM by 13.7% and 23.2%, respectively. In terms of Precision, our approach yields a 16.7% increase over the Anomaly Transformer and a 4.1% improvement over DCdetector. Although IForest shows a slightly higher precision of -1.5% difference compared to our model, this minor trade-off is balanced by significant gains in recall.
On the NIPS-TS-SWAN dataset, our model demonstrates exceptional performance. Notably, our model surpasses the Anomaly Transformer with improvements of 5.2% in precision, 12.5% in the recall, and 11.5% in the F1-score. Comparisons with DCdetector show more modest advancements, with gains of 0.4% in precision, 0.3% in recall, and 0.4% in the F1-score. In contrast, our enhancements compared to IForest are particularly significant, reflecting an absolute increase of 39.0% in precision and up to 56.9% and 15.5% in the F1-score. Additionally, when compared to OCSVM, our model reveals an improvement of 48.5% in precision, a 25.3% increase in the F1-score, and a 10.1% boost in recall. These results highlight the architectural strengths of our approach, particularly the Sparse Attention mechanism and Mamba-SSM, which adeptly identify subtle anomalies while minimizing noise interference. While minor trade-offs have been observed such as a slight reduction in precision on the NIPS-TS-GECCO dataset and a marginal decrease in recall on the NIPS-TS-SWAN dataset our model consistently balances precision and recall to achieve optimal overall performance.